OpenStudy (anonymous):

What are the approximate solutions of 2x2 − 7x = 3 rounded to the nearest hundredth? No real solutions x ≈ −0.77 and x ≈ 7.77 x ≈ −0.39 and x ≈ 3.89 x ≈ −3.89 and x ≈ 0.39

OpenStudy (ucreategames):

Im confused by this problem. 2x2 would make people think you're multiplying 2 by 2. What does the equation really look like?

OpenStudy (anonymous):

2x^2 - 7x = 3

OpenStudy (ucreategames):

Oh ok

OpenStudy (jack1):

factor first: (2x-1)(x-3)

OpenStudy (jack1):

so either 2x-1=0 or x-3 = 0

OpenStudy (jack1):

so x= 1/2 or x = 3

OpenStudy (anonymous):

that is not in the choices though

OpenStudy (jack1):

yeah... somethings wrong...

OpenStudy (jack1):

2x^2 - 7x = 3 2x^2 - 7x - 3 = 0 (2x-1)(x-3) ===> this line's wrong

OpenStudy (anonymous):

so how do i solve it?

OpenStudy (jack1):

y = ax^2 + bx + c y = 0 at x points of:

OpenStudy (jack1):

\[ \frac{ -b +/-\sqrt{(b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

i'm confused

OpenStudy (jack1):

so: from your equation a = 2 b = -7 c = -3 sub those numbers into the equation I gave you and you'll get answers for x

OpenStudy (jack1):

\[x = \frac{ 7 - \sqrt{73} }{ 4}\]

OpenStudy (jack1):

\[or x = \frac{ 7 + \sqrt{73} }{ 4}\]

OpenStudy (jack1):

so x = 3.886 and x = -0.386

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (jack1):

all good, sorry for the stuff up earlier

OpenStudy (anonymous):

thats fine